Jim had an appointment at 1:00 pm at a location 50 miles from his home. Driving at the speed limit of 30 mph , he arrived 20 minutes late. what time did Jim leave his house?
Answer by richwmiller(17219)

30*t=50
t=50/30
t=5/3
t=1 2/3
20 minutes =1/3
he left 1 1/3 hours before 1 pm
one hour before is noon
He left 20 minutes before noon
He left home at 11:40 am